# Wireless Power Transfer in Space using Flexible, Lightweight, Coherent Arrays

### 27 Jan 2024 v1

@@ 3.1 In-Orbit Wireless Power Transfer Experimental Results
The rectified voltage curves corresponding to optimizations performed in LEO are pictured in Fig. 8(b) and 8(a) [^4] The curves correspond to the optimizations that resulted in the highest rectified power to the respective rectennas during the duration of the mission.. The flex array created coherent focal points on each of the rectennas. The peak rectified DC powers received by the the sidewall and broadside rectennas were 231mW and 251mW, respectively. Each of the 16 stacks, composed of two PA cores (Fig. 4), consumed about 230mA from a 1.73V supply. Once phases were optimized to maximize power rectified by the receiver, power was switched from a resistive load to a boost circuit which powers an LED. The intensity of the LED, shown in Fig. 11, demonstrates the magnitude of power transferred.
@@ 3.2 Individual Cores: Phase Shifting Performance
where $\omega$ is the wave frequency, $\phi_{m}(c)$ is core $m$’s CMU phase function for code $c$, $n\in\{m-1,m+1\}$ is the index of the other core in the stack, $A_{m}$ is the amplitude of core $m$’s outgoing wave, and $P_{m}e^{j\phi_{m}^{P}}$ is the propagation vector from core $m$ to the receiver (with amplitude $P_{m}$ and phase $\phi_{m}^{P}$). The propagation vector accounts for an “infinite” sum of reflections emanating from the core which can vary owing to (mostly) perturbations on the MLI blanket and possibly other variable, internal physical features. This vector can vary, but is quasi-static for the measurement duration and likely changes negligibly in-orbit where there is no wind or (expected) physical change to the payload. The magnitude of the electric field can be expressed more simply as:
@@ 5.1.1 Rectified Voltage
Primary among the anomalous symptoms observed during in-orbit experiments are properties of the rectified voltage waveforms produced at the rectenna. During optimizations, rectified voltage should increase more or less monotonically, save for expected and occasional dramatically low dips in voltage as the algorithm stumbles on local maxima in the search space. After 10-30 seconds, the algorithm should stabilize at a maximum voltage (often close to the global maximum), which it is able to hold for the remainder of the optimization period. Inability to remain at this maximum, demonstrated by a slow drift downward (“drooping”) or sporadic changes to the output voltage, indicates the system is varying in time and undermining the previously computed search space optimum.
Some of the results measured in-orbit are anomalous with respect to both the absolute voltage level achieved by normal optimization runs and the shape of the rectified voltage curve. With respect to the absolute voltage level: after the first optimization, the system is unable to reproduce peak voltage at the receiver. Optimizations stabilize, instead, at lower voltage values.
With respect to the shape of the voltage curves: optimizations often failed to produce the typical increasing asymptotic curve that is indicative of a healthy, working system. Instead, curves showed sporadic behavior including dramatic but continuous decreases in rectified voltage and disjointed decreases in rectified voltage. This is evidenced by the data presented from the first in-orbit test performed on March 3rd, 2023 in Fig. 19(b).
Presented in Fig. 19(a) is data collected during integration testing wherein the expected voltage curves are produced by optimization. During a typical optimization, voltage levels repeatedly and consistently asymptotically increase and then stabilize at some maximum. When an optimization begins at the stable point of the previous optimization, as is done in the second half of the data in Fig. 19(a), there should be marginal increases in rectified voltage, but generally a flat curve.
@@ 5.1.2 PA Supply Current
Power amplifier (PA) supply currents measured during in-orbit tests also appeared d in some cases. PA supply lines are routed from the flight computer (FC) to the AP where they are voltage-regulated and then passed to the FA for routing to the two RFICs. Four PA supply lines (1A, 1B, 2A, 2B) route to the two RFICs (1/2); for RFIC1, quads 1 and 2 share supply line 1A and quads 3 and 4 share supply line 1B. On the AP, these lines are voltage- and current-monitored during regular “health checks.” These health checks provide the telemetry data referenced herein which are indicative of PA operation during testing.
@@ 5.1.3 Chip-Chip Interference Pattern
Figure 22: Rectified voltage during linear sweep of phase offset added to DB2.

### 16 Feb 2024 v2

@@ 3.1 In-Orbit Wireless Power Transfer Experimental Results
The rectified power curves corresponding to optimizations performed in LEO are pictured in Fig. 8(b) and 8(a) [^4] The curves correspond to the optimizations that resulted in the highest rectified power to the respective rectennas during the duration of the mission.. The flex array created coherent focal points on each of the rectennas. The peak rectified DC powers received by the the sidewall and broadside rectennas were 231mW and 251mW, respectively. Each of the 16 stacks, composed of two PA cores (Fig. 4), consumed about 230mA from a 1.73V supply. Once phases were optimized to maximize power rectified by the receiver, power was switched from a resistive load to a boost circuit which powers an LED. The intensity of the LED, shown in Fig. 11, demonstrates the magnitude of power transferred.
@@ 3.2 Individual Cores: Phase Shifting Performance
where $\omega$ is the wave frequency, $\phi_{m}(c)$ is core $m$’s CMU phase function for code $c$, $n\in\{m-1,m+1\}$ is the index of the other core in the stack, $A_{m}$ is the amplitude of core $m$’s outgoing wave, and $P_{m}e^{j\phi_{m}^{P}}$ is the propagation vector from core $m$ to the receiver (with amplitude $P_{m}$ and phase $\phi_{m}^{P}$). The propagation vector accounts for an “infinite” sum of reflections emanating from the core which can vary owing (mostly) to perturbations on the MLI blanket and possibly other variable, internal physical features. This vector can vary, but is quasi-static for the measurement duration and likely changes negligibly in-orbit where there is no wind or (expected) physical change to the payload. The magnitude of the electric field can be expressed more simply as:
@@ 5.1.1 Rectified Power
Primary among the anomalous symptoms observed during in-orbit experiments are properties of the rectified power waveforms produced at the rectenna. During optimizations, rectified power should increase more or less monotonically, save for expected and occasional dramatically low dips in power as the algorithm stumbles on local maxima in the search space. After 10-30 seconds, the algorithm should stabilize at a maximum power (often close to the global maximum), which it is able to hold for the remainder of the optimization period. Inability to remain at this maximum, demonstrated by a slow drift downward (“drooping”) or sporadic changes to the output power, indicates the system is varying in time and undermining the previously computed search space optimum.
Some of the results measured in-orbit are anomalous with respect to both the absolute power level achieved by normal optimization runs and the shape of the rectified power curve. With respect to the absolute power level: after the first optimization, the system is unable to reproduce peak power at the receiver. Optimizations stabilize, instead, at lower power values.
With respect to the shape of the power curves: optimizations often failed to produce the typical increasing asymptotic curve that is indicative of a healthy, working system. Instead, curves showed sporadic behavior including dramatic but continuous decreases in rectified power and disjointed decreases in rectified power. This is evidenced by the data presented from the first in-orbit test performed on March 3rd, 2023 in Fig. 19(b).
Presented in Fig. 19(a) is data collected during integration testing wherein the expected power curves are produced by optimization. During a typical optimization, power levels repeatedly and consistently asymptotically increase and then stabilize at some maximum. When an optimization begins at the stable point of the previous optimization, as is done in the second half of the data in Fig. 19(a), there should be marginal increases in rectified power, but generally a flat curve.
@@ 5.1.2 PA Supply Current
Power amplifier (PA) supply currents measured during in-orbit tests also appeared in some cases. PA supply lines are routed from the flight computer (FC) to the AP where they are voltage-regulated and then passed to the FA for routing to the two RFICs. Four PA supply lines (1A, 1B, 2A, 2B) route to the two RFICs (1/2); for RFIC1, quads 1 and 2 share supply line 1A and quads 3 and 4 share supply line 1B. On the AP, these lines are voltage- and current-monitored during regular “health checks.” These health checks provide the telemetry data referenced herein which are indicative of PA operation during testing.
@@ 5.1.3 Chip-Chip Interference Pattern
Figure 22: Rectified power during linear sweep of phase offset added to DB2.